# Evolutionary Algorithm with Detection Region Method for Constrained Multi-Objective Problems with Binary Constraints

### v1

@@ Abstract
Solving constrained multi-objective optimization problems (CMOPs) is a challenging task. While many practical algorithms have been developed to tackle CMOPs, real-world scenarios often present cases where the constraint functions are unknown or unquantifiable, resulting in only binary outcomes (feasible or infeasible). This limitation reduces the effectiveness of constraint violation guidance, which can negatively impact the performance of existing algorithms that rely on this approach. Such challenges are particularly detrimental for algorithms employing the $\varepsilon$-based method, as they hinder effective relaxation of the feasible region. To address these challenges, this paper proposes a novel algorithm called DRMCMO based on the detection region method. In DRMCMO, detection regions dynamic monitor feasible solutions to enhance convergence, helping the population escape local optima. Additionally, these regions collaborate with the neighbor pairing strategy to improve population diversity within narrow feasible areas. We have modified three existing test suites to serve as benchmark test problems for CMOPs with binary constraints(CMOP/BC) and conducted comprehensive comparative experiments with state-of-the-art algorithms on these test suites and real-world problems. The results demonstrate the strong competitiveness of DRMCMO against state-of-the-art algorithms. Given the limited research on CMOP/BC, our study offers a new perspective for advancing this field.
{IEEEImpStatement}
CMOPs are common in real-world optimization scenarios, with those involving binary constraints presenting significant challenges for existing algorithms. This work introduces DRMCMO, a novel algorithm designed to address CMOPs with binary constraints. The detection region method empowers traditional algorithms to solve these challenges effectively, and its high portability indicates that it may serve as a valuable tool for optimization researchers and practitioners in the future. Comprehensive experiments comparing DRMCMO with state-of-the-art algorithms demonstrate its competitive performance.
{IEEEkeywords}
@@ 1 Introduction
In real-world optimization, there exists a class of constraints that may be unknown or unquantifiable, introducing additional complexity. For example, collision avoidance is a critical constraint in autonomous driving that often provides only binary feedback [4]. Similarly, in the Job Shop Problem, strict precedence relationships exist between certain operations [5], yielding binary outcomes: if these constraints are not satisfied, the solution is deemed infeasible. In multi-robot path planning, constraints that prevent collisions between robots also provide only binary feedback, without quantifiable results [6].
These non-quantifiable constraint conditions are common in CMOPs and require careful consideration. Recently, Li et al. [7] defined a specific class of CMOPs with unknown constraints, which they call CMOP/UC. In fact, the CMOP/UC they consider is a CMOP with binary constraint conditions—constraints that can only return 0 (satisfied) or 1 (violated). This paper will refer to this problem as CMOP/BC, i.e., ”Constrained Multi-Objective Problems with Binary Constraints.”
@@
Over the past two decades, evolutionary algorithms(EAs) have significantly advanced in solving multiobjective optimization problems [8]. These multiobjective optimization algorithms(MOEAs) can be broadly categorized into three types: 1) Dominance-based algorithms: Examples include NSGA-II [9] and SPEA2 [10], which excel at problems with conflicting objectives and produce well-distributed Pareto-optimal solutions. 2) Decomposition-based algorithms: Algorithms such as MOEA/D [11] and MSOPS-II [12] are advantageous in many objectives problems, providing satisfactory Pareto sets with limited resources. 3) Indicator-based algorithms: Algorithms such as IBEA [13] and HypE [14] use performance indicators to guide the selection process towards well-distributed and high-quality solutions. Each algorithm type has its strengths and weaknesses, making it essential for researchers to select or hybridize them based on problem characteristics.
@@
However, these MOEAs cannot directly solve CMOPs as they lack essential components for handling constraints [15]. Therefore, various effective constraint-handling techniques (CHTs) have been developed over the past few decades to enable MOEAs to tackle CMOPs. These CHTs can be broadly categorized into the following approaches: 1) Feasibility-oriented methods [9] [16]: These methods prioritize feasible solutions during the selection process, which may result in excessive selection pressure on feasible solutions. 2) $\varepsilon$-based methods [17] [18]: These methods solve CMOPs by adjusting constraint boundaries, though selecting an appropriate $\varepsilon$ variant can be challenging. 3) Multi-objective optimization-based methods [19] [20]: Some methods reformulate the constrained problem as a multiobjective optimization problem, treating constraint violations as additional objectives, which can complicate objective optimization. 4) Stochastic ranking [21]: This approach probabilistically favors individuals with superior objective convergence while occasionally selecting those with better constraint violations. However, not all solutions exhibit dominance relationships, and determining these probabilities requires careful consideration. 5) Penalty functions [22] [23]: These algorithms incorporate constraint information into the fitness function, penalizing infeasible solutions, but tuning the penalty parameters can be challenging. These fundamental CHTs can be integrated into different MOEA frameworks. For example, multi-population or multi-stage constrained MOEAs may employ a combination of these techniques to effectively solve complex CMOPs. By leveraging these advanced CHTs, MOEAs have become increasingly capable of addressing various real-world optimization problems with diverse constraints.
@@
    Proposing the Detection Region Method (DRM): We introduce a detection region method to address the challenges of solving constrained multiobjective optimization problems with binary constraints (CMOP/BC), particularly the limitations of $\varepsilon$-based methods. This method dynamically relaxes the selection pressure on feasible solutions and monitors promising feasible solutions to assist the population in escaping local optima.
@@
The remainder of this paper is organized as follows. Section 2 presents the foundational concepts relevant to this study and reviews current constrained multiobjective evolutionary algorithms (CMOEAs). Section 3 details the proposed algorithm and its components. Sections 4 and 5 outline the experimental settings and analyze the results of a series of experiments, respectively. Finally, Section 6 concludes the paper and discusses potential directions for future work.
2 Related Work
2.1 Basic Definitions
2.1.1 CMOP
@@ 2.1.5 Pareto Dominance
$\ \begin{cases}\forall i\in(1,2,\ldots,m):f_{i}(\mathbf{x_{1}})\leq f_{i}(\mathbf{x_{2}})\\ \exists j\in(1,2,\ldots,m):f_{j}(\mathbf{x_{1}})<f_{j}(\mathbf{x_{2}})\end{cases}$ (6)
@@
2.1.6 Constraint Dominance Principle
@@
2.2 Constrained Multiobjective Evolutionary Algorithms
@@
2.2.1 Feasibility-Guided CMOEAs
@@
2.2.2 Constraint Relaxation-Based CMOEAs
@@
2.2.3 Multi-Population, Multi-Stage, or Multi-Task CMOEAs
@@
Remark 3: Algorithms that rely on constraint relaxation, particularly those using a dynamic threshold to relax CV, experience the most significant performance loss when addressing CMOP/BC. This is because the approach of relaxing CV becomes entirely ineffective for this class of problems. We will elaborate on this point in section 2 C
@@
2.3 Current Challenges in Existing Work for CMOP/BC
@@ 3 The Proposed Algorithm
3.1 Detection Region Method
@@
$\begin{cases}\textit{CV}(\mathbf{x})=0\\ \quad\quad\text{or}\\ \varphi(\mathbf{x})=\bigvee_{\textit{p}\in\textit{centers}}(||\mathbf{F}(\mathbf{x})-\textit{p}||<\textit{r})\end{cases}$ (8)
@@
2:     if $\mathbf{x}$ satisfies the conditions of Equation (8then
@@
7:  $\textit{Pop}_{\textit{k}}$leftarrow Select the N best solutions from Q through environmental selection;
@@ 3.2 Framework
1:  $\textit{Pop}_{0}$leftarrow Generate the initial population of size $N$;
@@
13:        end if
14:        $\alpha$leftarrow update by Equation (10);
15:        r $\leftarrow$ update by Equation (9);
16:        $\textit{Pop}_{\textit{k}}$leftarrow environmental selection for Q based on DRM (Algorithm 1);
17:     else
18:        $\textit{r}^{\textit{max}}$ $\leftarrow||\mathbf{F}^{\text{min}}||$;
@@
20:        $\textit{Pop}_{\textit{k}}$leftarrow environmental selection for Q based on CDP;
@@
[…]
@@
Algorithm 1 outlines the pseudocode for our framework. Lines 1 and 2 involve the initialization of the population and the archive, respectively. In line 3, $\textit{r}^{\textit{max}}$ denotes the maximum radius of the detection regions, initially set to the magnitude of the minimal objective vector of the population. Line 4 indicates k, the current generation, while line 5 records $\textit{k}_{\textit{s}}$, the generation when DRM is activated. Lines 7-8 detail the reproduction and merging processes of the populations. Lines 18 to 20 highlight that in the early stages of evolution, if no feasible solutions exist in the archive, the search for feasible solutions primarily relies on the CDP for constraint handling. Although the guiding role of CV diminishes when addressing constrained multiobjective optimization problems with CMOP/BC, CDP continues to exert significant selective pressure on feasible solutions. The calculation of fitness and environment selection are basically consistent with SPEA2 [10]. Meanwhile, $\textit{r}^{\textit{max}}$ is continuously updated during the convergence process. Once a feasible solution is found in the archive, the algorithm switches strategies to our proposed DRM (lines 14 to 16) to reduce the selective pressure of CDP on feasible solutions, thereby preventing convergence to local optima. DRM is conceptually similar to $\varepsilon$-based methods, both aiming to dynamically control the relaxation of constraints to guide the population’s search behavior. However, as discussed in Section 3 C, $\varepsilon$-based methods have critical shortcomings in dealing with CMOP/BC, which motivated the development of our DRM.
@@
3.3 Computational Complexity of DRMCMO
@@ 4 Experimental Setup
4.1 Benchmark Test Problems
@@
These modified CMOP test problems retain the original characteristics and are perfectly adapted to the binary constraint nature of CMOP/BC, thereby facilitating the evaluation of algorithms in dealing with CMOP/BC. The number of evaluations for all problems is set to 100000. For parameter settings (objective number M and decision variable number D) of MW/BC, see Table 1, and for LIRCMOP/BC and DASCMOP/BC, refer to Table 2.
4.2 Algorithm Settings
@@ 4.3 Performance Metric
5 Experimental Results and Discussion
5.1 Experimental Results on Benchmark Test Problems
Table 1: IGD Results Obtained by the Proposed DRMCMO and All the Compared CMOEAs on MW/BC Problems.
@@ 5.1 Experimental Results on Benchmark Test Problems
The MW test suite encompasses a diverse range of complex constrained scenarios, including small feasible regions, dispersed feasible regions, nonlinear constraints, and various shapes of CPF. Building upon the foundation of the original MW problems, the MW/BC introduces even more challenges. All constraints in this version do not provide precise, quantifiable feedback; they only provide binary outcomes. Experimental results have shown that this modification presents severe challenges to existing state-of-the-art CMOEAs. Table 1 illustrates the average IGD and standard deviation obtained by the proposed DRMCMO and other compared CMOEAs over 30 independent runs on the MW/BC test suite. DRMCMO achieved the best IGD on 22 questions. Compared with CMOEA-MS, MSCMO, MCCMO, cPDEA, EMCMO, and IMTCMO, DRMCMO is significantly better in 12, 10, 7, 9, and 9 problems, respectively. We conducted a Friedman test on the MW/BC suite. As displayed in Fig. 8, the results of the average friedman rankings for all algorithms indicate that DRMCMO leads with a ranking of 1.51, significantly outperforming its peers.
@@
Table 2: IGD Results Obtained by the Proposed DRMCMO and All the Compared CMOEAs on LIRCMOP/BC and DASCMOP/BC.
@@
LIRCMOP is mainly designed to test the ability of algorithms to cross large infeasible regions and also covers a variety of complex constraint scenarios. Based on LIRCMOP, the LIRCMOP/BC version binarizes the return values of all constraints; that is, the constraints will not give precise and quantifiable results but only provide binary results. This change increases the difficulty of algorithm processing. Table 2 lists the average IGD and standard deviation obtained by the proposed DRMCMO and other compared CMOEAs after 30 independent runs on the LIRCMOP/BC and DASCMOP/BC test suites. DRMCMO achieved the best IGD performance on 22 problems. Compared with CMOEA-MS, MSCMO, MCCMO, cPDEA, EMCMO, and IMTCMO, DRMCMO significantly outperformed the compared algorithms on 22, 19, 18, 20, 19 and 20 problems. Fig. 9 shows the results obtained by all algorithms on LIRCMOP/BC6. The results clearly show the significant gap between other algorithms and DRMCMO, among which only the final population of DRMCMO successfully covers the entire CPF. It is worth noting that both CMOEA-MS and IMTCMO are trapped in local optimality and fail to cross large infeasible regions effectively. This is mainly because when the constraints can only provide binary feedback, the effect of the fitness function that depends on the CV is significantly reduced, resulting in the performance of CMOEA-MS being limited. At the same time, $\varepsilon$-based methods relied on in IMTCMO is also completely ineffective, failing to properly relax the restrictions on the feasible region, which in turn causes the population to fall into local optimality. In the comparison of IGD convergence curves of all algorithms on LIRCMOP/BC5 in Fig. 7 (a), it can be seen that DRMCMO also has the fastest convergence speed. In order to further demonstrate the competitiveness and superior performance of DRMCMO, we performed a Friedman test on LIRCMOP/BC. In the average Friedman ranking results of all algorithms shown in Fig. 8, DRMCMO performed best with a ranking of 1.46 and significantly outperformed other algorithms.
@@
The DASCMOP test suite also covers a variety of complex constraint scenarios. Based on the original DASCMOP problem, we propose the DASCMOP problem with binary constraints (DASCMOP/BC). From the DASCMOP/BC experimental comparison data in Table 2, we can see that DRMCMO performs significantly better than other algorithms. Fig. 12 shows the populations obtained by all algorithms on the DASCMOP/BC3 problem, among which only DRMCMO successfully searched the complete CPF. Due to the weakening of CV guidance and the failure of $\varepsilon$-based methods, some algorithms such as CMOEA-MS, cDPEA, EMOCMO, and IMTCMO have fallen into local optima. Although MSCMO and MCCMO have roughly searched CPF, these two algorithms need to consume many evaluation times to maintain performance. Under the 100000 evaluation limit set in this paper, their performance is significantly inferior to the proposed DRMCMO. In the comparison of IGD convergence curves for all algorithms on DASCMOP/BC2, as shown in Fig. 7 (b), it is evident that DRMCMO exhibits the fastest convergence speed. Further analysis, judging from the average Friedman ranking results of all algorithms shown in Fig. 8, DRMCMO performed the best with a ranking of 1.22 and was significantly better than other algorithms. Comprehensive experimental results on these three CMOP/BC test problem sets have proven that compared with existing top CMOEAs, our proposed DRMCMO has higher adaptability and competitiveness in processing CMOP/BC.
@@
Table 3: HV Results Obtained by the Proposed DRMCMO and All the Compared CMOEAs on three Real-World problems.
@@
5.2 Ablation Study Results Analysis

### v2

# A Detection Region Method-Based Evolutionary Algorithm for Binary Constrained Multiobjective Optimization

@@ Abstract
Solving constrained multi-objective optimization problems is highly challenging, especially when constraint functions are unknown or unquantifiable, yielding only binary outcomes (feasible or infeasible). Such cases significantly limit the performance of many advanced constrained multi-objective evolutionary algorithms (CMOEAs), particularly those relying on $\varepsilon$-based methods. To address these challenges, this paper proposes a novel CMOEA framework based on detection region method (DRMCMO). In DRMCMO, detection regions dynamically monitor feasible solutions to enhance convergence. By using existing feasible solutions as a reference, the detection regions are dynamically adjusted to help the population escape local optima. We have modified three existing test suites to serve as benchmark problems for Constrained Multi-Objective Problems with Binary Constraints (CMOP/BC) and conducted comprehensive comparative experiments with state-of-the-art algorithms on these test suites and real-world problems. The results demonstrate that DRMCMO is highly competitive against other state-of-the-art algorithms. Given the limited research on CMOP/BC, our study provides a new perspective for advancing this field.
Index Terms:
@@ I Introduction
In practical optimization scenarios, a significant category of constraints often remains unknown or unquantifiable, thereby adding substantial complexity to the problem-solving process. For instance, in autonomous driving, collision avoidance is a crucial constraint that typically provides only binary feedback: either a collision is avoided (feasible) or it occurs (infeasible) [4]. Similarly, in the Job Shop Scheduling Problem, strict precedence relationships between operations must be maintained [5]; failure to satisfy these constraints renders the solution infeasible. In multi-robot path planning, constraints that prevent collisions between robots also offer only binary feedback, with no quantifiable measures of constraint violation [6]. Non-quantifiable constraints are prevalent in CMOPs and demand meticulous consideration. Recently, Li et al. [7] identified a subclass of CMOPs with unknown constraints, termed CMOP/UC. This subclass essentially involves binary constraints, where the constraints can only return two possible outcomes: 0 (satisfied) or 1 (violated). For clarity and consistency, this paper refers to this problem as CMOP/BC, short for ”Constrained Multiobjective optimization Problems with Binary Constraints.”
@@
Over the past two decades, evolutionary algorithms(EAs) have significantly advanced in solving multiobjective optimization problems [8]. These multiobjective optimization evolutionary algorithms(MOEAs) can be broadly categorized into three types: 1) Dominance-based algorithms: Examples include NSGA-II [9] and SPEA2 [10], which excel at problems with conflicting objectives and produce well-distributed Pareto-optimal solutions. 2) Decomposition-based algorithms: Algorithms such as MOEA/D [11] and MSOPS-II [12] are advantageous in many objectives problems, providing satisfactory Pareto sets with limited resources. 3) Indicator-based algorithms: Algorithms such as IBEA [13] and HypE [14] use performance indicators to guide the selection process towards well-distributed and high-quality solutions. Each algorithm type has its strengths and weaknesses, making it essential for researchers to select or hybridize them based on problem characteristics.
@@
However, these MOEAs cannot directly solve CMOPs as they lack essential components for handling constraints [15]. Therefore, various effective constraint-handling techniques (CHTs) have been developed over the past few decades to enable MOEAs to tackle CMOPs. These CHTs can be broadly categorized into the following approaches: 1) Feasibility-oriented methods [9] [16]: These methods prioritize feasible solutions during the selection process, which may result in excessive selection pressure on feasible solutions. 2) $\varepsilon$-based methods [17] [18]: These methods solve CMOPs by adjusting constraint boundaries, though selecting an appropriate $\varepsilon$ variant can be challenging. 3) Multi-objective optimization-based methods [19] [20]: Some methods reformulate the CMOP as a multiobjective optimization problem, treating constraint violations as additional objectives, which can complicate objective optimization. 4) Stochastic ranking [21]: This approach probabilistically favors individuals with superior objective convergence while occasionally selecting those with better constraint violations. However, not all solutions exhibit dominance relationships, and determining these probabilities requires careful consideration. 5) Penalty functions [22] [23]: These algorithms incorporate constraint information into the fitness function, penalizing infeasible solutions, but tuning the penalty parameters can be challenging. These fundamental CHTs can be integrated into different MOEA frameworks. For example, multi-population or multi-stage constrained MOEAs may employ a combination of these techniques to effectively solve complex CMOPs. By leveraging these advanced CHTs, MOEAs have become increasingly capable of addressing various real-world optimization problems with diverse constraints.
@@
    Proposing the Detection Region Method (DRM): We introduce a detection region method to address the challenges of solving CMOP/BC, particularly the limitations of $\varepsilon$-based methods. This method dynamically relaxes the selection pressure on feasible solutions and monitors promising feasible solutions to assist the population in escaping local optima.
@@
The remainder of this paper is organized as follows. Section II presents the foundational concepts relevant to this study and reviews current constrained multiobjective evolutionary algorithms (CMOEAs). Section III details the proposed algorithm and its components. Sections IV and V outline the experimental settings and analyze the results of a series of experiments, respectively. Finally, Section VI concludes the paper and discusses potential directions for future work.
II Related Work
II-A Basic Definitions
II-A1 CMOP
@@ II-A5 Pareto Dominance
[…]
@@
II-A6 Constraint Dominance Principle
@@
II-B Constrained Multiobjective Evolutionary Algorithms
@@
II-B1 Feasibility-Guided CMOEAs
@@
II-B2 Constraint Relaxation-Based CMOEAs
@@
II-B3 Multi-Population, Multi-Stage, or Multi-Task CMOEAs
@@
Remark 3: Algorithms that rely on constraint relaxation, particularly those using a dynamic threshold to relax CV, experience the most significant performance loss when addressing CMOP/BC. This is because the approach of relaxing CV becomes entirely ineffective for this class of problems. We will elaborate on this point in section II C
@@
II-C Current Challenges in Existing Work for CMOP/BC
@@ III The Proposed Algorithm
III-A Detection Region Method
@@
[…]
@@
2:  if $\mathbf{x}$ satisfies the conditions of Equation (8) then
@@
[…]
@@ III-B Framework
[…]
@@
12:    $\textit{k}_{\textit{s}}$ $\leftarrow\textit{k}$;
13:   end if
14:   $\alpha$ $\leftarrow$ update by Equation (10);
15:   r $\leftarrow$ update by Equation (9);
16:   $\textit{Pop}_{\textit{k}}$ $\leftarrow$ environmental selection for Q based on DRM (Algorithm 1);
17:  else
@@
20:   $\textit{Pop}_{\textit{k}}$ $\leftarrow$ environmental selection for Q based on CDP;
@@
22:  Arch $\leftarrow$ UpdateArchive($\textit{Pop}_{\textit{k}}\cupArch$) based on CDP;
@@
Algorithm 1 outlines the pseudocode for our framework. Lines 1 and 2 involve the initialization of the population and the archive, respectively. In line 3, $\textit{r}^{\textit{max}}$ denotes the maximum radius of the detection regions, initially set to the magnitude of the minimal objective vector of the population. Line 4 indicates k, the current generation, while line 5 records $\textit{k}_{\textit{s}}$, the generation when DRM is activated. Lines 7-8 detail the reproduction and merging processes of the populations. Lines 18 to 20 highlight that in the early stages of evolution, if no feasible solutions exist in the archive, the search for feasible solutions primarily relies on the CDP for constraint handling. Although the guiding role of CV diminishes when addressing CMOP/BC, CDP continues to exert significant selective pressure on feasible solutions. The calculation of fitness and environment selection are basically consistent with SPEA2 [10]. Given a solution $\mathbf{x}$ and a solution set $P$, the fitness can calculated as follow:
$fitness(\mathbf{x})=\sum_{\mathbf{y}\in P,\mathbf{y}\prec\mathbf{x}}|R^{\mathbf{x}}|+\frac{1}{\sigma_{\mathbf{x}}^{k}+2},$ (12)
where $R^{\mathbf{x}}$ measures convergence and represents the total number of solutions in the set $P$ that dominate $\mathbf{x}$ based on CDP. $\sigma_{\mathbf{x}}^{k}$ is the euclidean distance between $\mathbf{x}$ its $k$-th neighbor. Meanwhile, $\textit{r}^{\textit{max}}$ is continuously updated during the convergence process. Once a feasible solution is found in the archive, the algorithm switches strategies to our proposed DRM (lines 14 to 16) to reduce the selective pressure of CDP on feasible solutions, thereby preventing convergence to local optima. DRM is conceptually similar to $\varepsilon$-based methods, both aiming to dynamically control the relaxation of constraints to guide the population’s search behavior. However, as discussed in Section III C, $\varepsilon$-based methods have critical shortcomings in dealing with CMOP/BC, which motivated the development of our DRM.
@@
III-C Discussion
In comparing DRMCMO with $\varepsilon$-constrained relaxation algorithms, the discussion focuses on how the DRM effectively overcomes the fundamental limitations of $\varepsilon$-based methods when dealing with CMOP/BCs, thus demonstrating its unique advantages. $\varepsilon$-based constrained relaxation methods relax constraint violations by introducing a dynamic threshold, treating solutions with CV values below this threshold as ”pseudo-feasible” solutions. This approach aims to expand the feasible region and alleviate excessive selection pressure towards feasible solutions. However, as the paper highlights, in CMOP/BC problems, the binary nature of the constraint functions (returning only 0 or 1) makes CV quantification inaccurate, and CV values are often identical across many regions. This renders the $\varepsilon$-based methods ineffective, as its $\varepsilon$-feasible region” completely coincides with the true feasible region, failing to achieve effective constraint relaxation.
In contrast, DRM offers a novel and more adaptive approach. Instead of relying on inaccurate CV quantification, DRM relaxes constraints by establishing dynamic ”detection regions” in the objective space around discovered feasible solutions. The radii of these regions are dynamically adjusted throughout the evolutionary process, starting with a larger radius to help solutions trapped in local optima escape. As the algorithm converges, the detection regions gradually shrink, guiding the population to explore areas closer to the true constrained Pareto front and enhancing solution diversity. Furthermore, whenever a new, high-quality feasible solution is found, DRM uses it as a new reference point to expand its lateral search scope, creating a mutual reinforcement mechanism between feasible solutions and detection regions that promotes a more effective search process.
Therefore, the advantage of DRMCMO lies in its robust framework, specifically designed to handle the challenges posed by CMOP/BCs. It relaxes constraints through a method independent of traditional CV quantification, making it more effective than Epsilon methods that rely on CV information. This approach not only mitigates selection pressure but also enables a more efficient exploration and convergence towards the true feasible region.
III-D Computational Complexity of DRMCMO
@@ IV Experimental Setup
IV-A Benchmark Test Problems
@@
These modified CMOP test problems retain the original characteristics and are perfectly adapted to the binary constraint nature of CMOP/BC, thereby facilitating the evaluation of algorithms in dealing with CMOP/BC. The number of evaluations for all problems is set to 100000. For parameter settings (objective number M and decision variable number D) of MW/BC, see Table I, and for LIRCMOP/BC and DASCMOP/BC, refer to Table II.
IV-B Algorithm Settings
@@ IV-C Performance Metric
V Experimental Results and Discussion
V-A Experimental Results on Benchmark Test Problems
TABLE I: IGD Results Obtained by the Proposed DRMCMO and All the Compared CMOEAs on MW/BC Problems.
@@ V-A Experimental Results on Benchmark Test Problems
The MW test suite encompasses a diverse range of complex constrained scenarios, including small feasible regions, dispersed feasible regions, nonlinear constraints, and various shapes of CPF. Building upon the foundation of the original MW problems, the MW/BC introduces even more challenges. All constraints in this version do not provide precise, quantifiable feedback; they only provide binary outcomes. Experimental results have shown that this modification presents severe challenges to existing state-of-the-art CMOEAs. Table I illustrates the average IGD and standard deviation obtained by the proposed DRMCMO and other compared CMOEAs over 30 independent runs on the MW/BC test suite. DRMCMO achieved the best IGD on 22 questions. Compared with CMOEA-MS, MSCMO, MCCMO, cPDEA, EMCMO, and IMTCMO, DRMCMO is significantly better in 12, 10, 7, 9, and 9 problems, respectively. We conducted a Friedman test on the MW/BC suite. As displayed in Fig. 8, the results of the average friedman rankings for all algorithms indicate that DRMCMO leads with a ranking of 1.51, significantly outperforming its peers.
@@
TABLE II: IGD Results Obtained by the Proposed DRMCMO and All the Compared CMOEAs on LIRCMOP/BC and DASCMOP/BC.
@@
LIRCMOP is mainly designed to test the ability of algorithms to cross large infeasible regions and also covers a variety of complex constraint scenarios. Based on LIRCMOP, the LIRCMOP/BC version binarizes the return values of all constraints; that is, the constraints will not give precise and quantifiable results but only provide binary results. This change increases the difficulty of algorithm processing. Table II lists the average IGD and standard deviation obtained by the proposed DRMCMO and other compared CMOEAs after 30 independent runs on the LIRCMOP/BC and DASCMOP/BC test suites. DRMCMO achieved the best IGD performance on 22 problems. Compared with CMOEA-MS, MSCMO, MCCMO, cPDEA, EMCMO, and IMTCMO, DRMCMO significantly outperformed the compared algorithms on 22, 19, 18, 20, 19 and 20 problems. Fig. 9 shows the results obtained by all algorithms on LIRCMOP/BC6. The results clearly show the significant gap between other algorithms and DRMCMO, among which only the final population of DRMCMO successfully covers the entire CPF. It is worth noting that both CMOEA-MS and IMTCMO are trapped in local optimality and fail to cross large infeasible regions effectively. This is mainly because when the constraints can only provide binary feedback, the effect of the fitness function that depends on the CV is significantly reduced, resulting in the performance of CMOEA-MS being limited. At the same time, $\varepsilon$-based methods relied on in IMTCMO is also completely ineffective, failing to properly relax the restrictions on the feasible region, which in turn causes the population to fall into local optimality. In the comparison of IGD convergence curves of all algorithms on LIRCMOP/BC5 in Fig. 7 (a), it can be seen that DRMCMO also has the fastest convergence speed. In order to further demonstrate the competitiveness and superior performance of DRMCMO, we performed a Friedman test on LIRCMOP/BC. In the average Friedman ranking results of all algorithms shown in Fig. 8, DRMCMO performed best with a ranking of 1.46 and significantly outperformed other algorithms.
@@
The DASCMOP test suite also covers a variety of complex constraint scenarios. Based on the original DASCMOP problem, we propose the DASCMOP problem with binary constraints (DASCMOP/BC). From the DASCMOP/BC experimental comparison data in Table II, we can see that DRMCMO performs significantly better than other algorithms. Fig. 12 shows the populations obtained by all algorithms on the DASCMOP/BC3 problem, among which only DRMCMO successfully searched the complete CPF. Due to the weakening of CV guidance and the failure of $\varepsilon$-based methods, some algorithms such as CMOEA-MS, cDPEA, EMOCMO, and IMTCMO have fallen into local optima. Although MSCMO and MCCMO have roughly searched CPF, these two algorithms need to consume many evaluation times to maintain performance. Under the 100000 evaluation limit set in this paper, their performance is significantly inferior to the proposed DRMCMO. In the comparison of IGD convergence curves for all algorithms on DASCMOP/BC2, as shown in Fig. 7 (b), it is evident that DRMCMO exhibits the fastest convergence speed. Further analysis, judging from the average Friedman ranking results of all algorithms shown in Fig. 8, DRMCMO performed the best with a ranking of 1.22 and was significantly better than other algorithms. Comprehensive experimental results on these three CMOP/BC test problem sets have proven that compared with existing top CMOEAs, our proposed DRMCMO has higher adaptability and competitiveness in processing CMOP/BC.
@@
TABLE III: HV Results Obtained by the Proposed DRMCMO and All the Compared CMOEAs on three Real-World problems.
@@
V-B Ablation Study Results Analysis